# Vib2Mol: from vibrational spectra to molecular structures—a versatile deep learning model

### v3

@@ 2.4 Spectrum-structure correlation captured by Vib2Mol
From the macro perspective, the effectiveness of Vib2Mol was explored on the test set of VB-Mols-Raman. Figure 4C visualizes the distribution of cosine similarities between spectrum-structure embeddings of individual molecule before and after training by Vib2Mol. The similarity distribution before training was concentrated around -0.023, and significantly increased to 0.86 after training, demonstrating the effective alignment between spectral and structural representations of each molecule upon training with Vib2Mol. To further investigate if our model truly understanding of the spectrum-structure relationship, we analyze its performance on failure cases. Figure S1 illustrates the Tanimoto similarity distributions between incorrectly generated or retrieved molecules and their targets. Among the incorrectly generated results, approximately 56% of the molecules had a Tanimoto similarity greater than 0.5 to their target molecule (0.5 is a common threshold for determining molecular similarity[14, 21]), with the overall distribution having an expectation of 0.53. Among the incorrectly retrieved molecules, about 47% had a Tanimoto similarity greater than 0.5, and the expectation for this distribution is 0.47. These results suggest that even when the outputs were incorrect, they still maintained a degree of structural similarity to their intended targets, such as misplacing a methyl group or an oxygen atom.
@@ 2.5 Generating products for chemical reaction
The autonomous robot laboratory is leading new paradigm shifts in fields such as chemical synthesis, catalysis and drug screening[1]. The related advancements are underpinned by spectral information provided by various techniques. For instance, the autonomous and efficient exploration of chemical synthesis (such as combinatorial small-molecule synthesis, designing supramolecular materials, and screening photocatalysts) can be achieved with the aid of HPLC-MS and NMR[2]. However, applying current spectrum-to-structure methods to real conditions remains challenging. On the one hand, researchers have different levels of knowledge about different synthetic methods. As a result, it is crucial to fully utilize the available prior knowledge to help select appropriate molecular elucidation strategies. On the other hand, spectra measured in practice are often mixtures of reactants and products. It is a key issue to elucidate molecules under the interference of impurities. The solvation of these two problems by Vib2Mol were demonstrated as follows.
@@
To better reflect real-world conditions, we evaluated Vib2Mol in interpreting mixture spectra of products and reactants. By extracting approximately 15,000 chemical reactions listed in the second World AI4S Prize-Material Science Track[22], we calculated the Raman spectra of reactants and products, and mixed them according to the yields (ignoring the differences in Raman scattering cross-sections of different molecules) to simulate the mixed spectra measured in real condition (see Supplementary Information for details). Spectrum-structure retrieval and de novo generation were used to simulate the scenarios where the expected product is either present in or absent from the database, respectively.
@@ 2.6 Peptide sequencing and PTMs identification
Native proteins are composed of 20 amino acids and their post-translational modifications (PTMs). As sequences determine the structures and functions of proteins, protein sequencing and the identification of PTMs sites are key issues in reveals the functions and mechanisms of proteins in cellular function regulation, gene expression regulation, signal transduction, and the occurrence and development of diseases. To simplify the complexity of protein sequences, a bottom-up strategy is commonly adopted, which involves generating peptides of varying lengths (1-4 amino acids or longer) through chemical or enzymatic cleavage. By sequentially identifying these peptides, de novo sequencing can be achieved. However, considering the vast sequence space of polypeptides ($20^{N}_{AA}$), efficiently identifying the 20 amino acids and their combinations remains highly challenging[23]. Although the unique fingerprint vibrational information in Raman spectrum is for each biomolecule (i.e., DNA, proteins)[24, 25, 26], the complexity of Raman spectra of peptides hindered the systematic identification of polypeptides for de novo protein sequencing. Limiting the length of peptide sequences to tetrapeptides or shorter, we tried to infer peptide sequence using its Raman spectra by Vib2Mol.
The Vib2Mol pre-trained on VB-Mols was fine-tuned by peptides represented by SMILES. As shown in Table 1, the Recall@1 of the model (Vib2Mol-SMILES) for spectrum-structure retrieval is 62.97%, when the to-be-determined peptide is in the database. Otherwise, the Recall@1 for de novo generation drops to 50.07%. The low Recall@1 ignited us to change peptide representation from SMILES to residue sequences, considering the relatively patterned residue structure. The obtained Vib2Mol-sequence model significantly improved Recall@1 for retrieval and generation up to 70.33% and 51.49%, respectively. This improvement is mainly because of the drastically reduced token length by residue sequences, thereby reducing the complexity and improving the accuracy of sequence generation. This is the reason why current models[3, 12, 27] are mainly based on residue sequences. It is not surprising that the Vib2Mol sequence performed best in elucidating dipeptides (97.37% and 86.84% for retrieval and generation, respectively). Since the search space increases exponentially with the number of residues, the performance of Vib2Mol-sequence gradually decreases with the increasing length of peptides. Nevertheless, even for tetrapeptides, the Recall@1 for retrieval and generation can still reach 68.03% and 48.18%, respectively.
@@ 3 Discussion
Vib2Mol also demonstrates the potential for in situ monitoring of dynamic chemical reactions and life processes on the basis of vibrational spectroscopy. In the future, to better elucidate molecular conformations in dynamic processes, a possible improvement lies in the introduction of stereochemical information. In addition, it is also of great interest to design more flexible generative modules to equip the models with bidirectional spectrum-to-structure and structure-to-spectrum predictions.
@@ 4.1 Reference data
Density functional theory (DFT) was employed to perform conformational optimization of molecules and calculated the corresponding infrared and Raman spectra in VB. Unless otherwise specified, all quantum chemical calculations were carried out using the Gaussian 09 program[28]. The geometries were optimized using the B3LYP-D3BJ functional with a 6-311+G** basis set. Frequency calculations were obtained at the same level at the optimized geometry.
For details on the specific division of the above datasets for training, validation, and testing, please refer to Table S6. To facilitate subsequent calculations, the spectral dimensions were unified to 1024, and molecular structures were all represented using SMILES.
@@ 4.2 Related works and baseline models
To fairly compare the performance among Vib2Mol and current methods, we surveyed spectrum-to-structure models based on vibrational spectroscopy in the community. For spectral-structure retrieval, CL is currently the most popular framework. DeepSearch[3], CReSS[5], SMEN[29] and vibraCLIP[19] all employ CL to bring the spectra and corresponding structures of the same molecule closer together, achieving great spectral-structure retrieval performance for mass spectrometry, NMR, IR, and Raman respectively. We fully followed the training and inference configurations of SMEN and vibraCLIP to test their performance on the aforementioned benchmarks. It is worth noting that SMEN, vibraCLIP, and Vib2Mol represent molecules with coordinates (conformation), molecular graphs, and SMILES, respectively. Conformation offers a significant advantage in theoretical spectrum-structure retrieval because a precisely corresponding conformation provides a better molecular representation. However, accurate conformations are unavailable for experimental benchmarks, making only conformation-independent molecular representations easily transferable.
For conditional generation, MLM is currently the most popular approach. CO-BERT realized bidirectional prediction between molecular structures and vibrational spectra by MLM[30]. However, CO-BERT focus on predicting atomic coordinates by vibrational spectra and contextual structural information. Therefore, we built a similar variant based on BERT and compared it with Vib2Mol in Supplementary Information.
For de novo generation, both IR2Mol[14] and the Patch-based Self-Attention (PBSA) model[9] employ an encoder-decoder architecture and integrate molecular formula constraints. We followed the established training and inference protocols of IR2Mol and PBSA to assess their performance across the benchmarks. It is important to note that both of these baselines employ sophisticated data augmentation strategies, and PBSA additionally uses an ensemble learning approach. While we acknowledge the efficacy of these strategies, for the sake of fairness and efficiency in this study, we exclusively used data augmentation applied only to SMILES strings and did not incorporate any ensemble learning strategies in this paper. Recognizing that molecular formulae are not always easily obtainable[7], we developed two distinct versions of Vib2Mol—one that incorporates chemical formulae and another that operates without them. Both versions exhibited strong performance.
@@ 4.4 Alignment between spectrum and molecular structure
During the training phase, we used a symmetric cross-entropy loss[31] to calculate the similarity errors between the spectra and structures of the same molecule and updated the neural network based on this. The specific formula of the loss function is as follows:
$L_{total}=\frac{1}{2}(L_{spectrum}+L_{structure})=-\frac{1}{2}(\sum_{i=1}^{m}{ilog(p_{i})}+\sum_{j=1}^{n}{jlog(q_{j})})$ (1)
where $m$ and $n$ are the number of rows and columns of the probability distribution matrix, respectively. $ilog(p_{i})$andj $ljlog(q_{j})$represent the cross-entropy of spectrum-to-structure, and structure-to-spectrum, respectively.
During the testing or inference phase, only the dot product of the features of the to-be-determined spectrum and the molecules in the library needs to be calculated, and the top-k results are taken as the final results (Figure S4).
@@ 4.5 Spectrum-structure matching and re-ranking
Building on a large-scale spectrum-structure alignment achieved through contrastive learning, we found that a small number of negative pairs—those with similar spectra or molecular structures—were still difficult to distinguish. To address this, we introduced the spectrum-structure matching. This module is a binary classification task, where the model uses a linear layer to predict whether a spectrum-structure pair is positive (matched) or negative (unmatched) given their multimodal feature. In order to find more informative negatives, we adopt the hard negative mining strategy by BLIP[32], where negatives pairs with higher contrastive similarity in a batch are more likely to be selected to compute the loss. During inference, this module naturally scores the retrieved or generated molecular structures against a to-be-determined spectrum, enabling a deep learning-based re-ranking of candidate molecules.

### v4

@@ 2.5 Spectrum-structure correlation captured by Vib2Mol
From the macro perspective, the effectiveness of Vib2Mol was explored on the test set of VB-Mols-Raman. Figure 4C visualizes the distribution of cosine similarities between spectrum-structure embeddings of individual molecule before and after training by Vib2Mol. The similarity distribution before training was concentrated around -0.023, and significantly increased to 0.86 after training, demonstrating the effective alignment between spectral and structural representations of each molecule upon training with Vib2Mol. To further investigate if our model truly understanding of the spectrum-structure relationship, we analyze its performance on failure cases. Figure S1 illustrates the Tanimoto similarity distributions between incorrectly generated or retrieved molecules and their targets. Among the incorrectly generated results, approximately 56% of the molecules had a Tanimoto similarity greater than 0.5 to their target molecule (0.5 is a common threshold for determining molecular similarity[alberts2024leveraging, hu2025deep]), with the overall distribution having an expectation of 0.53. Among the incorrectly retrieved molecules, about 47% had a Tanimoto similarity greater than 0.5, and the expectation for this distribution is 0.47. These results suggest that even when the outputs were incorrect, they still maintained a degree of structural similarity to their intended targets, such as misplacing a methyl group or an oxygen atom.
@@ 2.6 Generating products for chemical reaction
The autonomous robot laboratory is leading new paradigm shifts in fields such as chemical synthesis, catalysis and drug screening[burger2020mobile]. The related advancements are underpinned by spectral information provided by various techniques. For instance, the autonomous and efficient exploration of chemical synthesis (such as combinatorial small-molecule synthesis, designing supramolecular materials, and screening photocatalysts) can be achieved with the aid of HPLC-MS and NMR[dai2024autonomous]. However, applying current spectrum-to-structure methods to real conditions remains challenging. On the one hand, researchers have different levels of knowledge about different synthetic methods. As a result, it is crucial to fully utilize the available prior knowledge to help select appropriate molecular elucidation strategies. On the other hand, spectra measured in practice are often mixtures of reactants and products. It is a key issue to elucidate molecules under the interference of impurities. The solvation of these two problems by Vib2Mol were demonstrated as follows.
@@
To better reflect real-world conditions, we evaluated Vib2Mol in interpreting mixture spectra of products and reactants. By extracting approximately 15,000 chemical reactions listed in the second World AI4S Prize-Material Science Track[sais2024], we calculated the Raman spectra of reactants and products, and mixed them according to the yields (ignoring the differences in Raman scattering cross-sections of different molecules) to simulate the mixed spectra measured in real condition (see Supplementary Information for details). Spectrum-structure retrieval and de novo generation were used to simulate the scenarios where the expected product is either present in or absent from the database, respectively.
@@ 2.7 Peptide sequencing and PTMs identification
Native proteins are composed of 20 amino acids and their post-translational modifications (PTMs). As sequences determine the structures and functions of proteins, protein sequencing and the identification of PTMs sites are key issues in reveals the functions and mechanisms of proteins in cellular function regulation, gene expression regulation, signal transduction, and the occurrence and development of diseases. To simplify the complexity of protein sequences, a bottom-up strategy is commonly adopted, which involves generating peptides of varying lengths (1-4 amino acids or longer) through chemical or enzymatic cleavage. By sequentially identifying these peptides, de novo sequencing can be achieved. However, considering the vast sequence space of polypeptides ($20^{N}_{AA}$), efficiently identifying the 20 amino acids and their combinations remains highly challenging[pappas2016dynamic]. Although the unique fingerprint vibrational information in Raman spectrum is for each biomolecule (i.e., DNA, proteins)[chen2018high, zhao2022label, li2022enhanced], the complexity of Raman spectra of peptides hindered the systematic identification of polypeptides for de novo protein sequencing. Limiting the length of peptide sequences to tetrapeptides or shorter, we tried to infer peptide sequence using its Raman spectra by Vib2Mol.
The Vib2Mol pre-trained on VB-Mols was fine-tuned by peptides represented by SMILES. As shown in Table 1, the Recall@1 of the model (Vib2Mol-SMILES) for spectrum-structure retrieval is 62.97%, when the to-be-determined peptide is in the database. Otherwise, the Recall@1 for de novo generation drops to 50.07%. The low Recall@1 ignited us to change peptide representation from SMILES to residue sequences, considering the relatively patterned residue structure. The obtained Vib2Mol-sequence model significantly improved Recall@1 for retrieval and generation up to 70.33% and 51.49%, respectively. This improvement is mainly because of the drastically reduced token length by residue sequences, thereby reducing the complexity and improving the accuracy of sequence generation. This is the reason why current models[yu2025towards, mao2023mitigating, tran2019deep] are mainly based on residue sequences. It is not surprising that the Vib2Mol sequence performed best in elucidating dipeptides (97.37% and 86.84% for retrieval and generation, respectively). Since the search space increases exponentially with the number of residues, the performance of Vib2Mol-sequence gradually decreases with the increasing length of peptides. Nevertheless, even for tetrapeptides, the Recall@1 for retrieval and generation can still reach 68.03% and 48.18%, respectively.
@@
To rigorously evaluate the generalization capability of the model, we benchmarked Vib2Mol against a home-made experimental Raman spectral dataset of peptides, comprising 20 amino acids, 14 dipeptides, 13 tripeptides, and 2 tetrapeptides (Figure S7). A 5-fold cross-validation on the 29 multi-residue peptide samples (excluding monomeric amino acids) was conducted, with the performance metrics summarized in Table 3.
In the retrieval task, Vib2Mol achieved an average Recall@1 of 72.00% and Recall@3 of 86.66%, demonstrating its proficiency in mapping noisy experimental signals to correct structural identities. Notably, retrieval Recall scaled positively with peptide length. This is primarily due to the distribution of our dataset within the latent space. As quantified in Table S9, dipeptides exhibit higher average cosine similarity. This indicates that shorter sequences reside in a “crowded” region of the representation space, where a higher density of neighboring negative samples leads to retrieval ambiguity. Conversely, longer peptides act as “informative outliers” in sparser regions, allowing for more precise identification despite their inherent spectral complexity.
In the de novo generation, the Avg. Recall@1 of13.34% highlights the significant gap between theoretically simulated dataset and experimental one, let alone the extreme scarcity of labeled experimental Raman spectra. Further, a declined performance is for de novo generation as the peptide chain elongated. This inverse correlation is anticipated, as the chemical search space expands exponentially with each additional amino acid residue, substantially increasing the complexity of sequence reconstruction. While de novo generation on complex experimental samples remains a frontier challenge, these results provide empirical evidence of the Vib2Mol’s retrieval capabilities within real-world biological scenarios.
@@
Table 3: Benchmarking Vib2Mol on experimental peptide Raman spectra
Peptide Retrieval De novo generation
Recall@1 Recall@3 Recall@1 Recall@3
Dipeptide 64.28 78.57 28.57 28.57
Tripeptide 76.92 92.31 0 7.69
Tetrapeptide 100 100 0 0
5-fold avg. 72.00±9.56 86.66±12.47 13.34±12.47 16.66±18.25
@@ 3 Discussion
Vib2Mol also demonstrates the potential for in situ monitoring of dynamic chemical reactions and life processes using vibrational spectroscopy, particularly in discovery-oriented and automated experimental settings, where reaction pathways, intermediates, or products cannot be predefined a priori. In the future, to better elucidate molecular conformations in dynamic processes, a possible improvement lies in the introduction of stereochemical information. Furthermore, the current framework relies heavily on a SMILES-based molecular encoder, which restricts its applicability to non-molecular systems such as polymers and minerals. Transitioning toward more flexible molecular representations, such as graph-based or 3D-aware encoders, would broaden the application of this method. Finally, there is significant potential in developing bidirectional generative modules that enable seamless transitions between spectrum-to-structure and structure-to-spectrum predictions, creating a truly reversible mapping for molecular design and analysis.
@@ 4.1 Reference data
Density functional theory (DFT) was employed to perform conformational optimization of molecules and calculated the corresponding infrared and Raman spectra in VB. Unless otherwise specified, all quantum chemical calculations were carried out using the Gaussian 09 program[g09]. The geometries were optimized using the B3LYP-D3BJ functional with a 6-311+G** basis set. Frequency calculations were obtained at the same level at the optimized geometry.
Our datasets were partitioned based on unique identifiers (GDB9-ID for QM9S and a combination of GDB9-ID/ZINC-ID for VB-Mols, etc.). We employed a stratified random sampling approach based on these IDs to construct the training, validation, and test sets. Detailed statistics regarding the specific dataset divisions are provided in Table S10. To facilitate subsequent calculations, the spectral dimensions were unified to 1024, and molecular structures were all represented using SMILES.
@@ 4.2 Related works and baseline models
To fairly compare the performance among Vib2Mol and current methods, we surveyed spectrum-to-structure models based on vibrational spectroscopy in the community. For spectral-structure retrieval, CL is currently the most popular framework. DeepSearch[yu2025towards], CReSS[yang2021cross], SMEN[kanakala2024spectra] and vibraCLIP[rocabert2025multi] all employ CL to bring the spectra and corresponding structures of the same molecule closer together, achieving great spectral-structure retrieval performance for mass spectrometry, NMR, IR, and Raman respectively. We fully followed the training and inference configurations of SMEN and vibraCLIP to test their performance on the aforementioned benchmarks. It is worth noting that SMEN, vibraCLIP, and Vib2Mol represent molecules with coordinates (conformation), molecular graphs, and SMILES, respectively. Conformation offers a significant advantage in theoretical spectrum-structure retrieval because a precisely corresponding conformation provides a better molecular representation. However, accurate conformations are unavailable for experimental benchmarks, making only conformation-independent molecular representations easily transferable.
For conditional generation, MLM is currently the most popular approach. CO-BERT realized bidirectional prediction between molecular structures and vibrational spectra by MLM[yang2024cross]. However, CO-BERT focus on predicting atomic coordinates by vibrational spectra and contextual structural information. Therefore, we built a similar variant based on BERT and compared it with Vib2Mol in Supplementary Information.
For de novo generation, both IR2Mol[alberts2024leveraging] and the Patch-based Self-Attention (PBSA) model[wu2025transformer] employ an encoder-decoder architecture and integrate molecular formula constraints. We followed the established training and inference protocols of IR2Mol and PBSA to assess their performance across the benchmarks. It is important to note that both of these baselines employ sophisticated data augmentation strategies, and PBSA additionally uses an ensemble learning approach. While we acknowledge the efficacy of these strategies, for the sake of fairness and efficiency in this study, we exclusively used data augmentation applied only to SMILES strings and did not incorporate any ensemble learning strategies in this paper. Recognizing that molecular formulae are not always easily obtainable[hu2024accurate], we developed two distinct versions of Vib2Mol—one that incorporates chemical formulae and another that operates without them. Both versions exhibited strong performance.
@@ 4.4 Alignment between spectrum and molecular structure
During the training phase, we used a symmetric cross-entropy loss[radford2021learning] to calculate the similarity errors between the spectra and structures of the same molecule and updated the neural network based on this. The specific formula of the loss function is as follows:
$L_{total}=\frac{1}{2}(L_{spectrum}+L_{structure})=-\frac{1}{2B}(\sum_{i=1}^{B}{log(p_{i})}+\sum_{j=1}^{B}{log(q_{j})})$ (1)
Expanding this with the similarity scores, the formulation becomes:
$L_{total}=-\frac{1}{2B}(\sum_{i=1}^{B}{log(\frac{exp(s_{i,i}/\tau)}{\sum_{j=1}^{B}{exp(s_{i,j}/\tau)}})}+\sum_{j=1}^{B}{log(\frac{exp(s_{j,j}/\tau)}{\sum_{i=1}^{B}{exp(s_{i,j}/\tau)}})})$ (2)
where $B$ denotes the size of mini-batch, $p_{i}$ and $q_{j}$ represent the predicted probabilities of correct alignment of spectrum-to-structure, and structure-to-spectrum. These probabilities are calculated based on the similarity scores $s_{i,j}$ and scaled by a temperature hyperparameter $\tau$.
During the testing or inference phase, only the dot product of the features of the to-be-determined spectrum and the molecules in the library needs to be calculated, and the top-k results are taken as the results (Figure S4).
@@ 4.5 Spectrum-structure matching and re-ranking
Building on a large-scale spectrum-structure alignment achieved through contrastive learning, we found that a small number of negative pairs—those with similar spectra or molecular structures—were still difficult to distinguish. To address this, we introduced the spectrum-structure matching. This module is a binary classification task, where the model uses a linear layer to predict whether a spectrum-structure pair is positive (matched) or negative (unmatched) given their multimodal feature. In order to find more informative negatives, we adopt the hard negative mining strategy by BLIP[li2022blip], where negatives pairs with higher contrastive similarity in a batch are more likely to be selected to compute the loss. During inference, this module naturally scores the retrieved or generated molecular structures against a to-be-determined spectrum, enabling a deep learning-based re-ranking of candidate molecules.
The matching loss is formulated as a binary cross-entropy objective:
$L_{matching}=-\sum_{k\in\{pos,neg\}}y_{k}\log{\left({\hat{y}}_{k}\right)}$ (3)
where $y_{k}$ denotes the ground-truth binary label and $\hat{y}_{k}$ represents the predicted probability of a match.
@@ 4.7 Beam Search
4.8 Multimodal integration and mask modality modeling
The multimodal variant, Vib2Mol-MM, is designed to handle heterogeneous spectral data flexibly. The spectral input is structured as a tensor ${X}\in\operatorname{R}^{B\times C\times L}$, where $B$ is the batch size, $C$ denotes the number of spectral modalities (channels), $L$ represents the spectral dimension (the number of sampling points, which is 1024 in this work). We use a 1D convolution with an equal kernel size and stride to convert the input tensor into patch tokens. whose feature dimension is $D$ which is 768 in our case.
Raman and infrared spectra share a common x-axis (wavenumber), allowing them to be treated as complementary feature channels—analogous to the RGB channels in image recognition. This shared dimensionality enables a seamless transition between modalities by simply adjusting the input channel parameter, $C$, without requiring structural architectural changes. For unimodal input, $C=1$; for multimodal input, $C=2$, as the Raman ($S_{Raman}$) and infrared ($S_{IR}$) spectra are concatenated along the channel dimension. From a machine learning perspective, this approach allows the model to learn cross-modal correlations at each specific vibrational frequency. However, it is important to acknowledge that simple channel-wise concatenation is a relatively primitive approach to multimodal integration. It may not fully account for the Rule of Mutual Exclusion or the fundamental differences in selection rules—specifically, that IR transitions depend on changes in dipole moment, whereas Raman transitions depend on changes in polarizability.
To enable a unified framework capable of processing Raman-only, IR-only, or joint inputs, we propose a masked modality modeling training strategy. As detailed in Algorithm 1, we treat the two modalities as channels. During training, we randomly mask one modality with a probability $p_{mask}=0.5$, replacing the signal with a learnable embedding $E_{mask}$. This forces the model to learn robust inter-modal dependencies when both are present, while maintaining high performance when only a single modality is available.
Algorithm 1 Training Forward with Masked Modality Modeling
1:Raman spectra $S_{Raman}$, infrared spectra $S_{IR}$, Training flag $T$, Masking probability $p_{mask}$
2:Latent spectral representations $Z$
[…]
4:if $T$ is True and $\text{Rand}(0,1)<p_{mask}$ then
5:  $m\leftarrow\text{Sample}(\{0,1\})$ $\triangleright$ Randomly pick a modality to mask
6:  $X[:,m,:]\leftarrow E_{mask}$
7:end if
[…]
11:return $Z$
4.9 Model complexity and computational cost
We have summarized the model size (parameter count), computational complexity (GFLOPS), and memory requirements (VRAM) in Table S11. As illustrated, Vib2Mol achieves superior efficiency through extensive parameter sharing. For instance, in retrieval tasks, the integrated matching module and molecular encoder reduce the active parameter count relative to independent architectures. Similarly, for generation tasks, our shared decoder framework significantly lowers memory overhead while maintaining robust multi-task capabilities. Note that the VRAM requirements for training are based on the hyperparameters specified in Table S1-S2.
Furthermore, we conducted a comparative benchmark of inference times using an NVIDIA RTX 4090 GPU and an Intel Xeon Platinum 8468 CPU, as summarized in Table S12. The inference pipeline consists of three primary stages: pre-retrieval/generation, filtering, and re-ranking. While the pre-retrieval/ generation and re-ranking stages are optimized for GPU acceleration, the filtering stage is executed primarily on the CPU. Our results demonstrate that while the GPU offers superior throughput, the model remains highly performant on a CPU, with retrieval tasks consistently completing in under 100 ms.
4.10 Raman spectra acquisition and data processing
All the samples were placed on a glass slide wrapped with aluminum foil and gently pressed with a second glass slide to form a uniform and flat layer. Raman spectra were then obtained on Invia confocal Raman microscope (Renishaw, UK) with a 785 nm excitation laser. The laser power was set to 16.9 mW, with an integration time of 10 s per acquisition, and each spectrum was obtained by averaging three accumulations. Savitzky–Golay filter with a window size of 7 and an order of 3 was used to denoise. After the subtraction of the biofluorescence background, spectra were individually normalized with the maximum peak value.